# SmartShift: A Secure and Efficient Approach to Smart Contract Migration

### v1

@@ Abstract.
copyright: nonebooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayconference: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering; June 2025; Trondheim, Norwayccs: Computing methodologiesccs: Computing methodologies Distributed computing methodologies
@@ 1. Introduction
We structure the remainder of this paper as follows: in Section 2, we provide a critical review of related literature. In Section 3, we outline the formal foundation of SmartShift, detailing its design principles. In Section 4, we present the system architecture, detailing its core components. In Section 5, we evaluate SmartShift’s performance, demonstrating improved efficiency and reduced downtime. In Section 6, we analyze its security, emphasizing robustness against migration vulnerabilities. We conclude the paper in Section 7 by summarizing the key findings.
@@ 3. Formal Foundation
SmartShift optimizes migration by breaking down data transfers into manageable chunks and aligning function availability with data readiness, maintaining continuous operation during upgrades. The system uses a dependency matrix $\Delta$ and function priority vector $P$ to orchestrate migrations effectively. $\Delta$ maps function-data dependencies
$\Delta_{f_{i}\times d_{j}}=\begin{cases}1&\text{if function }f_{i}\text{ depends on data }d_{j}\\ 0&\text{otherwise}\end{cases}$
@@ 4. System Architecture
The architecture of SmartShift comprises interconnected components that intelligently analyze smart contract dependencies and orchestrate segmented migrations to optimize operational continuity. This section examines how these components interact within the migration workflow shown in Fig. 1, illustrating their collaborative roles in optimizing the contract transition process while maintaining operational continuity.
@@ 4.3.2. State Shard Generator
The State Shard Generator, precisely segments the smart contract’s state into individual shards, leveraging the detailed instructions provided by the State Extractor. Each shard encapsulates a single state variable or data element in the smart contract. This process involves selecting the appropriate slots from the contract’s storage, as defined by the State Extractor’s output, and partitioning them into discrete units or shards. Each shard is thus a self-contained packet of slots related to a state variable. This approach allows for targeted migrations where specific portions of the contract’s state can be updated or moved independently of others.
4.4. Batch Manager
The Batch Manager is a critical component of SmartShift, responsible for organizing and executing the migration process in a structured and efficient manner. As illustrated in Fig. 1 it consists of two subcomponents: the Batch Generator and the Batch Processor.
4.4.1. Batch Generator
We have developed the Batch Generator to sort the shards generated by the State Shard Generator using the dependency matrix and the priority vector. This ensures that interdependencies and operational importance are respected throughout the process. It adjusts the composition of shards as needed, either merging multiple shards together or breaking them down further. These adjustments are guided by gas limit constraints and user preferences. This flexibility allows us to create optimally sized batches that balance performance and resource usage while maintaining the integrity of the migration process.
4.4.2. Batch Processor
We encapsulate each batch into a blockchain transaction and sequentially send them to the blockchain using the Batch Processor. We have implemented It to process the batches in the order determined by their priority, ensuring that critical functions and data are migrated first. By encapsulating batches into transactions, the Batch Processor ensures that each migration step is executed securely and efficiently.
@@ 5. Implementation & Evaluation
To measure the performance of the solution, we have implemented a proof of concept for SmartShift. We implemented SmartShift’s key components in a combination of Python and Go. To access transaction history and gas limit required to implement the Priority Vector Calculator and the Batch Processor we have used services like Infura and Etherscan. Using this proof of concept, we analyzed how effectively our framework reduces migration-related downtime. For quantitative evaluation, we developed a metric called the Function Activation Threshold (FAT), which measures the average number of data elements required to activate smart contract functions. Lower thresholds indicate quicker migrations, while higher ones suggest longer unavailability. We analyzed this metric across the ERC-20, ERC-721, and ERC-1155 Ethereum token standards, assessing their impact on operations due to their critical role and widespread use in decentralized applications.
@@ 7. Conclusion
In this paper, we introduced SmartShift, a sophisticated framework that addresses key challenges related to smart contract migration security, efficiency, and scalability by employing granular state decomposition and incremental function activation. Our empirical tests demonstrate that SmartShift significantly reduces system downtime and strengthens defenses against common migration vulnerabilities. Through these innovations, SmartShift defines the next step in blockchain development, focusing on secure, seamless transitions for smart contracts

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728517isbn: 979-8-4007-1276-0/2025/06ccs: Computing methodologiesccs: Computing methodologies Distributed computing methodologies
@@ 1. Introduction
The rest of the paper is organized as follows: Section 2 reviews related literature, Section 3 introduces the design principles of SmartShift, Section 4 details its architecture, Section 5 evaluates performance, Section 6 analyzes security, and Section 7 concludes the paper.
@@ 3. Formal Foundation
SmartShift optimizes migration by breaking down data transfers into manageable chunks and aligning function availability with data readiness, maintaining continuous operation during upgrades. The system uses a dependency matrix $\Delta$ and function priority vector $P$ to orchestrate migrations effectively. $\Delta$ maps function-data dependencies $\Delta_{f_{i}\times d_{j}}=1$ if function $f_{i}$ depends on data $d_{j}$; otherwise, $\Delta_{f_{i}\times d_{j}}=0$.
@@ 4. System Architecture
The SmartShift architecture consists of interconnected components that analyze smart contract dependencies and coordinate segmented migrations to ensure continuity. This section explains their interaction within the migration workflow (Fig. 1).
@@ 4.3.2. State Shard Generator
The State Shard Generator precisely segments the smart contract’s state into individual shards, leveraging the detailed instructions provided by the State Extractor. Each shard encapsulates a single state variable or data element in the smart contract. This process involves selecting the appropriate slots from the contract’s storage, as defined by the State Extractor’s output, and partitioning them into discrete units or shards. Each shard is thus a self-contained packet of slots related to a state variable. This approach allows for targeted migrations where specific portions of the contract’s state can be updated or moved independently of others.
Batch Manager. The Batch Manager is a critical component of SmartShift, responsible for organizing and executing the migration process in a structured and efficient manner. As illustrated in Fig. 1 it consists of two subcomponents: the Batch Generator and the Batch Processor.
Batch Generator. We have developed the Batch Generator to sort the shards generated by the State Shard Generator using the dependency matrix and the priority vector. This ensures that interdependencies and operational importance are respected throughout the process. It adjusts the composition of shards as needed, either merging multiple shards together or breaking them down further. These adjustments are guided by gas limit constraints and user preferences. This flexibility allows us to create optimally sized batches that balance performance and resource usage while maintaining the integrity of the migration process.
Batch Processor. We encapsulate each batch into a blockchain transaction and sequentially send them to the blockchain using the Batch Processor. We have implemented It to process the batches in the order determined by their priority, ensuring that critical functions and data are migrated first. By encapsulating batches into transactions, the Batch Processor ensures that each migration step is executed securely and efficiently.
@@ 5. Implementation & Evaluation
To measure the performance of the solution, we have implemented a proof of concept for SmartShift. We implemented SmartShift’s key components in a combination of Python and Go. To access transaction history and gas limit required to implement the Priority Vector Calculator and the Batch Processor, we have used services like Infura and Etherscan. Using this proof of concept, we analyzed how effectively our framework reduces migration-related downtime. For quantitative evaluation, we developed a metric called the Function Activation Threshold (FAT), which measures the average number of data elements required to activate smart contract functions. Lower thresholds indicate quicker migrations, while higher ones suggest longer unavailability. We analyzed this metric across the ERC-20, ERC-721, and ERC-1155 Ethereum token standards, assessing their impact on operations due to their critical role and widespread use in decentralized applications.
@@ 7. Conclusion
In this paper, we introduced SmartShift, a sophisticated framework that addresses key challenges related to smart contract migration security, efficiency, and scalability by employing granular state decomposition and incremental function activation. Our empirical tests demonstrate that SmartShift significantly reduces system downtime and strengthens defenses against common migration vulnerabilities. Through these innovations, SmartShift defines the next step in blockchain development, focusing on secure, seamless transitions for smart contracts.